# The Cosmological Impact of Brane-Chern-Simons Massive Gravity

### v1

@@ 3 Perturbation Analysis
$S^{(2)}_{\rm B-C-S}=\frac{M_{\rm Pl}^{2}}{8}\int d^{3}k\,dt\,a^{3}N\Bigg{\{}\big{[}1+\varepsilon V(\varphi)\big{]}\bigg{[}\frac{\dot{h}_{ij}\dot{h}^{ij}}{N^{2}}-(\frac{k^{2}}{a^{2}}+\frac{4\dot{H}}{N}+6H^{2})h^{ij}h_{ij}\bigg{]}$
$+2\kappa V(\varphi)h^{ij}h_{ij}\Bigg{\}}.$
@@
$S^{(2)}_{\rm massive}=\frac{M_{\rm Pl}^{2}}{8}\int d^{3}k\,dt\,a^{3}Nm_{g}^{2}\Bigg{\{}\bigg{[}\alpha_{3}+\alpha_{4}-3\big{(}1+2\alpha_{3}+\alpha_{4}\big{)}a+2\big{(}3+3\alpha_{3}+\alpha_{4}\big{)}a^{2}\bigg{]}\frac{\dot{f}(t)}{a^{2}N}$
$-\bigg{[}1+2\alpha_{3}+\alpha_{4}-3\big{(}3+3\alpha_{3}+\alpha_{4}\big{)}a+2\big{(}6+4\alpha_{3}+\alpha_{4}\big{)}a^{2}\bigg{]}\frac{1}{a^{2}}\Bigg{\}}h^{ij}h_{ij}.$
@@
$M^{2}_{\rm GW}=\big{(}4\frac{\dot{H}}{N}+6H^{2}\big{)}\big{[}1+\varepsilon V(\varphi)\big{]}-2\kappa V(\varphi)+\Omega,$
@@
$\Omega=\Bigg{\{}2Na_{\rm SA}^{3\pm}\big{[}\alpha_{3}+3\big{]}-N\big{(}\alpha_{3}+1\big{)}+a_{\rm SA}^{2\pm}\bigg{[}\big{(}\alpha_{3}+2\big{)}\dot{f}(t)-2N\big{(}3\alpha_{3}+7\big{)}\bigg{]}$
$+a_{\rm SA}^{\pm}\bigg{[}N\big{(}5\alpha_{3}+8\big{)}-\dot{f}(t)\big{(}\alpha_{3}+1\big{)}\bigg{]}\Bigg{\}}\times\bigg{(}Na_{\rm SA}^{2\pm}\big{[}a_{\rm SA}^{2\pm}-1\big{]}\bigg{)}^{-1}.$ (28)
@@
It is important to note that we eliminate $\alpha_{4}$ using Eq. (16). The sign of the mass square of gravitational waves plays a crucial role in ensuring the stability of long-wavelength gravitational waves. When the mass square is positive, it indicates a stable configuration. Conversely, a negative mass square implies the presence of tachyons, leading to potential instability. However, due to the characteristic timescale associated with the tachyonic mass, which is on the order of the Hubble scale, the development of any instability would take a duration comparable to the age of the universe, thus having minimal impact on the stability of the system over observable timescales.
4 Cosmological Tests
The exploration of Type Ia supernovae provided compelling evidence for the accelerated expansion of the universe Copeland:2006wr ; Frieman:2008sn ; Perlmutter:2003kf ; Yang:2019fjt . Here, we utilize the Union2 supernovae Ia dataset, comprising 557 SNIa events Amanullah:2010vv , to investigate the Brane-Chern-Simons Massive Gravity theory. The observational data from the SNIa dataset will be presented in terms $\mu_{\rm obs}$, enabling a direct comparison with the theoretical predictions of the model under scrutiny.
$\mu_{\rm th}(z_{i})=5\log_{10}D_{L}(z_{i})+\mu_{0},$ (29)
it is important to note that $\mu_{0}=42.38-5\log_{10}\mathcal{H}$, where $\mathcal{H}$ represents the Hubble constant $H_{0}$ in units of $100\,{\rm km/s/Mpc}$. This relationship is essential for establishing the connection between the observed magnitude and the luminosity distance, which can be expressed as follows:
$D_{L}(z)=(1+z)\int_{0}^{z}\frac{dx}{E(n;q)},$ (30)
furthermore, it is evident that $E=H/H_{0}$ and $q$ are the model parameters. It is important to note that $X^{2}$ is defined as follows:
$X_{\mu}^{2}(q)=\sum_{i}\frac{[\mu_{\rm obs}(z_{i})-\mu_{\rm th}(z_{i})]^{2}}{\sigma^{2}(z_{i})},$ (31)
here, $\sigma$ corresponds to the $1\sigma$ error, and the parameter $\mu_{0}$ is a nuisance parameter independent of the data points. To minimize $X_{\mu}^{2}$ in equation (31), we expand it with respect $\mu_{0}$ Nesseris:2005ur ; DiPietro:2002cz .
$X_{\mu}^{2}(q)=\tilde{T}-2\mu_{0}\tilde{Y}+\mu_{0}^{2}\tilde{Z},$ (32)
where
$\tilde{T}(q)=\sum_{i}\frac{[\mu_{\rm obs}(z_{i})-\mu_{\rm th}(z_{i};\mu_{0}=0,q)]^{2}}{\sigma_{\mu_{\rm obs}}^{2}(z_{i})},$
$\tilde{Y}(q)=\sum_{i}\frac{\mu_{\rm obs}(z_{i})-\mu_{\rm th}(z_{i};\mu_{0}=0,q)}{\sigma_{\mu_{\rm obs}}^{2}(z_{i})},$
$\tilde{Z}=\sum_{i}\frac{1}{\sigma_{\mu_{\rm obs}}^{2}(z_{i})}.$ (33)
It should be explained that for $\mu_{0}=\frac{\tilde{Y}}{\tilde{Z}}$, equation (32) has a minimum at
$\tilde{X}_{\mu}^{2}(q)=\tilde{T}(q)-\frac{\tilde{Y}^{2}(q)}{\tilde{Z}}.$ (34)
As it is clear that $X_{\mu,{\rm min}}^{2}=\tilde{X}_{\mu,{\rm min}}^{2}$ we can consider minimizing $\tilde{X}_{\mu}^{2}$, which is independent of $\mu_{0}$. It is important to note that the best-fit model parameter is determined by minimizing $X^{2}=\tilde{X}_{\mu}^{2}$. At the same time, we know that the corresponding $\mathcal{H}$ is dependent on $\mu_{0}=\frac{\tilde{Y}}{\tilde{Z}}$ for the best-fit parameter. By considering equation (17) and applying the change of variables $a=\frac{1}{1+z}$ and $\frac{d}{dt}=-H(z+1)\frac{d}{dz}$, we can express the dimensionless Hubble parameter for this case. Consequently, the solution to the asymptotic state at small redshifts should be provided as follows:
$H(z)\sim H_{0}+\frac{\Theta^{\pm}}{H_{0}^{3}}z-\frac{3\Theta^{\pm\rm 2}+\Theta^{\pm}H_{0}^{4}}{2H_{0}^{7}}z^{2}+.....,$
where $\Theta^{\pm}$ is defined as
$\Theta^{\pm}=-\frac{2}{3}\frac{\Lambda_{\rm SA}^{\pm}+\kappa V(\varphi)}{\big{(}1+\varepsilon V(\varphi)\big{)}},$ (36)
it is important to note that $H_{0}$ represents the Hubble parameter at the present time. Therefore, the dimensionless Hubble parameter can be expressed as follows:
$E=\frac{H(z)}{H_{0}}=1+Fz-(\frac{3F^{2}+F}{2})z^{2}+......$
where
$F=\frac{\Theta^{\pm}}{H_{0}^{4}}.$ (38)
In this step, we present plots of $X^{2}$ and likelihoods as functions of the parameter $F$. Based on our calculations, the best-fit yields a value of $X_{\rm min}^{2}=544.044$, and the corresponding best-fit parameter is,
$F=$ $0.5497^{+0.0251}_{-0.0255},\quad\quad\mbox{with 1$\sigma$ uncertainty},$ (39)
$F=$ $0.5497^{+0.0500}_{-0.0515},\quad\quad\mbox{with 2$\sigma$ uncertainty}.$ (40)
Consequently, it is noteworthy to mention that the best-fit value for the Brane-Chern-Simons massive gravity theory corresponds to $\mathcal{H}=0.6986$.
Figure 1: The distance modulus diagram for the best fit (red solid line) of the parameter of the Brane-Chern-Simons massive gravity theory in comparison with the 557 Union2 SNIa data points (black dots).
Figure 2: The $X^{2}$ functions of parameter $F$ for the Brane-Chern-Simons massive gravity theory.
Figure 3: The likelihood as functions of parameter $F$ for the Brane-Chern-Simons massive gravity theory.
According to Figures 1 and 3, the result of fitting the Brane-Chern-Simons massive gravity theory to the cosmological data provides us with the optimal value for the model parameter.
Figure 4: The best-fit diagram for the luminosity distance $D_{L}$ with respect to velocity $V$ for the Brane-Chern-Simons massive gravity theory.
To evaluate the validity of this theory, we utilize the Union2 supernovae Ia dataset, consisting of $557$ SNIa events. By employing the Bayesian statistics method, we constrain the model parameters based on the SNIa data. Specifically, we generate plots of $X^{2}$ and the likelihood function as functions of the parameter $F$, allowing us to determine the best-fit value by minimizing $X^{2}$.
Our analysis yields a best-fit value of $F=0.5497$ corresponding to a minimum $X^{2}$ value of $544.044$. Using this optimal parameter value, we calculate the best-fit value of $\mathcal{H}$ for the theory, obtaining $\mathcal{H}=0.6986$.
Figure 1 illustrates the distance modulus diagram, depicting the relationship between the distance modulus $\mu$ and the redshift $z$, along with the best-fit curve for the Brane-Chern-Simons massive gravity theory based on supernovae Ia data. Additionally, we provide a plot of velocity $V$ versus luminosity distance $D_{L}$, demonstrating the agreement between theoretical predictions and supernovae Ia observations. In Figure 4, we present a plot of the velocity $V$ as a function of luminosity distance $D_{L}$ for the Brane-Chern-Simons massive gravity theory. The velocity is calculated from the redshift using the equation $V=c-\frac{2c}{(z+1)^{2}+1}$.
@@ 5 Conclusion
In this work, we introduce a novel extension of massive gravity theory, the Brane-Chern-Simons massive gravity theory. We explore its potential to explain the late-time acceleration of the universe by deriving the full set of equations of motion for an FLRW background and analyzing self-accelerating background solutions. A key finding is the identification of self-accelerating solutions characterized by an effective cosmological constant, $\Lambda_{\rm SA}^{\pm}$. These solutions exhibit well-behaved behavior without strong coupling issues, offering a promising framework to explain the accelerated expansion of the universe.
@@
The validation of our model is demonstrated in the final section through a comparison with the latest Union2 SNIa dataset, consisting of 557 supernovae events. The Brane-Chern-Simons massive gravity theory exhibits an excellent fit with the observational data, as illustrated in Fig. 1. This allows us to constrain the model parameters and obtain a best-fit value of $\mathcal{H}=0.6986$ for the Hubble parameter, reinforcing the theory’s capability to describe the late-time acceleration of the universe. Future studies, utilizing larger and more diverse datasets, will be crucial for further refining and validating these findings, ultimately advancing our understanding of the fundamental principles governing the universe’s behavior.
@@ Acknowledgements.
This work is based upon research funded by the University of Tabriz, Iran National Science Foundation (INSF), and Iran National Elites Foundation (INEF), under project No. 4014244. The authors thank A. Emir Gumrukcuolu for insightful comments. We also appreciate Nishant Agarwal for sharing notes and computational codes related to tensor perturbations, which greatly facilitated our research.

### v2

@@ 3 Perturbation Analysis
$S^{(2)}_{\rm B-C-S}=\frac{M_{\rm Pl}^{2}}{8}\int d^{3}k\,dt\,a^{3}N\Bigg{\{}\big{[}1+\varepsilon V(\varphi)\big{]}\bigg{[}\frac{\dot{h}_{ij}\dot{h}^{ij}}{N^{2}}-|K|(\frac{k^{2}}{a^{2}}+\frac{4\dot{H}}{N}+6H^{2})h^{ij}h_{ij}\bigg{]}$
$-\frac{2}{M_{\rm Pl}^{2}}\mathcal{L}_{m}+2\kappa V(\varphi)h^{ij}h_{ij}\Bigg{\}}.$
@@
$S^{(2)}_{\rm massive}=-\frac{M_{\rm Pl}^{2}}{4}\int d^{3}k\,dt\,m_{g}^{2}(a-\sqrt{|K|}f)\Bigg{\{}|K|f^{2}\bigg{[}\big{(}\alpha_{3}+\alpha_{4}\big{)}N-\alpha_{4}\dot{f}\bigg{]}+a^{2}\bigg{[}\big{(}6+4\alpha_{3}$
$+\alpha_{4}\big{)}N-\big{(}3+3\alpha_{3}+\alpha_{4}\big{)}\dot{f}\bigg{]}+\sqrt{|K|}af\bigg{[}\big{(}3\alpha_{3}+2\alpha_{4}\big{)}\dot{f}-\big{(}3+5\alpha_{3}+2\alpha_{4}\big{)}N\bigg{]}\Bigg{\}}h^{ij}h_{ij}.$
@@
$M^{2}_{\rm GW}=\big{(}4\frac{\dot{H}}{N}+6H^{2}\big{)}\big{[}1+\varepsilon V(\varphi)\big{]}|K|+\frac{2}{M_{\rm Pl}^{2}}p_{m}-2\kappa V(\varphi)+\Upsilon,$
@@
$\Upsilon=$ $\frac{m_{g}^{2}}{|K|^{4}(\alpha_{3}+\alpha_{4})^{4}\sqrt{|K|\zeta a}}\times\bigg{[}-|K|(\alpha_{3}+\alpha_{4})+\sqrt{|K|}(1+2\alpha_{3}+\alpha_{4})a\pm\sqrt{|K|\zeta a}\bigg{]}\Bigg{\{}|K|^{3}(\alpha_{3}$ (36)
$+\alpha_{4})^{3}(6+4\alpha_{3}+\alpha_{4})\sqrt{|K|\zeta a}\mp 2H|K|^{2}(\alpha_{3}+\alpha_{4})^{2}(3+3\alpha_{3}+\alpha_{4})a^{2}\bigg{(}\sqrt{|K|}(1+\alpha_{3}$
$+\alpha_{3}^{2}-\alpha_{4})a\pm(1+2\alpha_{3}+\alpha_{4})\sqrt{|K|\zeta a}\bigg{)}+\bigg{(}\pm\sqrt{|K|}(1+2\alpha_{3}+\alpha_{4})a+\sqrt{|K|\zeta a}\bigg{)}^{2}$
×$\times\bigg{(}|K|(\alpha_{3}+\alpha_{4})^{2}\sqrt{|K|\zeta a}-2H\alpha_{4}a^{2}\big{(}\pm\sqrt{|K|}\zeta+(1+2\alpha_{3}+\alpha_{4})\sqrt{|K|\zeta a}\big{)}\bigg{)}$
$-|K|(\alpha_{3}+\alpha_{4})\bigg{(}\sqrt{|K|}(1+2\alpha_{3}+\alpha_{4})a\pm\sqrt{|K|\zeta a}\bigg{)}\times\bigg{(}|K|(\alpha_{3}+\alpha_{4})(3$
$+5\alpha_{3}+2\alpha_{4})\sqrt{|K|\zeta a}\mp 2H(3\alpha_{3}+2\alpha_{4})a^{2}\big{[}\sqrt{|K|}\zeta$
±$\pm(1+2\alpha_{3}+\alpha_{4})\sqrt{|K|\zeta a}\big{]}\bigg{)}\Bigg{\}},\qquad where\qquad\zeta=a(1+\alpha_{3}+\alpha_{3}^{2}-\alpha_{4}).$
@@
The equation (36) is derived from substituting the solutions of $f$ from the equation (17). In Figures (1) and (2), we show the possible ranges of Eq. (36) which is the dispersion relation of gravitational wave originating from the massive gravity part. The sign of the mass square of gravitational waves plays a crucial role in ensuring the stability of long-wavelength gravitational waves. When the mass square is positive, it indicates a stable configuration. Conversely, a negative mass square implies the presence of tachyons, leading to potential instability. However, due to the characteristic timescale associated with the tachyonic mass, which is on the order of the Hubble scale, the development of any instability would take a duration comparable to the age of the Universe, thus having minimal impact on the stability of the system over observable timescales.
Figure 1: The possible ranges of $\gamma$, scale factor $a$, and $\alpha_{4}$ are determined by considering the values for $|K|=1$, $m_{g}=10^{-23}eV$ LIGOScientific:2017bnn , $H=69(km/s)/Mpc$, and $\alpha_{3}=5$.
Figure 2: The possible ranges of $\gamma$, scale factor $a$, and $\alpha_{3}$ are determined by considering the values for $|K|=1$, $m_{g}=10^{-23}eV$, $H=69(km/s)/Mpc$, and $\alpha_{4}=-5$.
The calculation of the dispersion relation for gravitational waves in the Brane-Chern-Simons massive gravity theory provides a representation of the propagation of gravitational perturbations in the Friedmann-Lemaître-Robertson-Walker (FLRW) cosmology. In principle, this propagation can be tested through observations of cosmological events, particularly those involving gravitational waves. The modifications introduced by this theory will result in additional contributions to the phase evolution of gravitational waveforms, which can be detected using precise matched-filtering techniques in data analysis Will:1997bb ; Mirshekari:2011yq .
Following the groundbreaking discovery of gravitational waves from a merging binary black hole (GW150914) by the LIGO/Virgo Collaboration, interest in testing the mass of the graviton has been rekindled LIGOScientific:2016lio ; LIGOScientific:2019fpa ; LIGOScientific:2020tif ; Shao:2020shv . Recent constraints on the graviton mass, derived from the combined analysis of gravitational wave events from the first and second transient catalogs, yield an upper limit of $m_{g}\leq 1.76\times 10^{-23}\,\mathrm{eV}/c^{2}$ at 90% credibility LIGOScientific:2020tif . Although the corresponding Compton wavelength is still much smaller than the Hubble scale, limiting its relevance to modified cosmology, ongoing and future efforts will continue to test this fundamental aspect of gravitation with an increasing number of gravitational wave events at various wavelengths. Notably, future space-based gravitational wave detectors will offer enhanced sensitivity to the graviton mass Will:1997bb .
In a related study, Nishizawa and Arai investigated the modified propagation of gravitational waves in a cosmological context using a parametrized framework, which accounted for a running Planck mass, a modified speed of gravitation, and anisotropic source terms Nishizawa:2017nef ; Arai:2017hxj ; Nishizawa:2019rra . Our results, representing a theory-specific analysis within the Brane-Chern-Simons massive gravity theory, complement their work and provide a foundation for future detailed analysis.
@@ 4 Conclusion
In this work, we introduce a novel extension of massive gravity theory, the Brane-Chern-Simons massive gravity theory. We explore its potential to explain the late-time acceleration of the Universe by deriving the full set of equations of motion for an FLRW background and analyzing self-accelerating background solutions. A key finding is the identification of self-accelerating solutions characterized by an effective cosmological constant, $\tilde{\rho}_{g}$. These solutions exhibit well-behaved behavior without strong coupling issues, offering a promising framework to explain the accelerated expansion of the Universe.
@@
The calculation of the dispersion relation in Brane-Chern-Simons massive gravity theory offers a means to test gravitational perturbation propagation in FLRW cosmology. Our analysis provides a foundation for future studies, complementing existing work on modified gravitational wave propagation and potentially shedding light on the nature of gravity.
@@ Acknowledgements.
This work is based upon research funded by the University of Tabriz, Iran National Science Foundation (INSF), and Iran National Elites Foundation (INEF), under project No. 4014244. This work is supported by the National Key R&\&&D Program of China (grant 2023YFE0117200) and the National Natural Science Foundation of China (Nos. 12105013). The authors thank A. Emir Gumrukcuolu for insightful comments. We also appreciate Nishant Agarwal for sharing notes and computational codes related to tensor perturbations, which greatly facilitated our research.